# Frequency-Time Diffusion with Neural Cellular Automata

### v1

@@ Introduction
Figure 3: Qualitative comparison between FourierDiff-NCA, VNCA, and DDM based on UNet M. With a parameter count of 886k, 9.73m, and 3.94m, respectively.
Data and Infrastructure
@@
CelebA dataset: The CelebA dataset (Liu et al. 2015), a widely used benchmark, consists of 202,599 images, each of size $178\times 218$. We scale all images to a uniform size of $64\times 64$ to match the dataset with the input requirements of the UNet and VNCA. The data is split 80%:10%:10% for training, validation, and testing. CelebA presents a challenge as it includes various facial images against different backgrounds. The inclusion of individuals with a range of accessories, hairstyles, and facial expressions further increases the complexity of the dataset. The size and variety of this dataset allow us to evaluate the ability of our models to handle intricate details and different visual features.
@@
BCSS Pathology dataset: This dataset (Amgad et al. 2019) contains 144 high-resolution pathology samples, which present a unique challenge for generative models. Initially, some images contain blur at the lowest level and are resized by a factor of four in x and y direction to obtain clear and concise visual patterns. We then extract patches of size $64\times 64$ for training purposes. The BCSS pathology dataset provides an opportunity to rigorously test the capabilities of our proposed methods in generating large images in the medical domain. For this dataset, we use an 80%:10%:10% split for training, validation, and testing.
@@
Infrastructure: All models are implemented in PyTorch (Paszke et al. 2019), where we use the official implementation of VNCA (Palm et al. 2022). The models are trained on an Ubuntu 22.04 system using an Nvidia RTX 3090 and an Intel i7-12700 processor. Additional details can be found in the supplementary material and the codebase that we will make available upon acceptance.
Metrics
@@
To assess the quality of the synthesized outputs, we use the well-established metrics Fréchet Inception Distance (FID) (Heusel et al. 2017) and Kernel Inception Distance (KID) (Bińkowski et al. 2018), which measure the similarity between the real and synthetic images based on the Inception-v3 (Szegedy et al. 2015) model. For both evaluations, we use a set of 2048 real images from the test split compared to an equal number of synthesized images.
Qualitative Comparison: Image Synthesis
@@
Examining the images, FourierDiff-NCA generated in Figure 3 shows that the model learns global information. Interestingly, FourierDiff-NCA does not have enough steps in the image space to achieve global knowledge transfer. Nevertheless, it exhibits such behavior, indicating that communication occurs in the Fourier space. Thanks to this Fourier communication, the model can capture various features of the underlying distribution, such as variations in facial features, hair color, clothing, and facial expressions, all while using a comparably small parameter count of 887k. The efficiency becomes apparent when comparing it to the ten times bigger VNCA and the five times bigger UNet M-based DDM, where the results look worse with blurry results and substantial color shifts. Although the results of FourierDiff-NCA are imperfect, they represent a promising step toward detailed image generation with Neural Cellular Automata and demonstrate the potential for complex generative tasks.
Figure 4: Influence of parameter count to image generation performance for FourierDiff-NCA, UNet, and VNCA (the detailed numbers can be found in the supplementary).
Quantitative Comparison: Image Synthesis
@@
Figure 4 compares FourierDiff-NCA, various UNet configurations, and VNCA. Notably, FourierDiff-NCA excels at the tradeoff between parameters and FID, outperforming larger UNet-based DDMs and VNCA. With 887k parameters, FourierDiff-NCA outperforms UNet-based DDMs with five times more parameters, along with VNCA, which has ten times the number of parameters. Although a large UNet with 111 million parameters produces the best results, performance drops with fewer parameters. Reducing the number of parameters in UNet configurations leads to noticeable degradation in FID and KID metrics. The smallest UNet configuration records an FID of 175.3 and a KID of 0.142. In contrast, FourierDiff-NCA, with 887k parameters, achieves an FID of 77.84 and a KID of 0.066. FourierDiff-NCA’s localized interaction architecture contributes to stable performance and improved parameter efficiency.
@@
Figure 5: Out-Of-Distribution image synthesis with FourierDiff-NCA of different scales and shapes.
Ablation
@@
The ablation study shown in Table 1 provides insights into the specific effects of the different hyperparameters of the FourierDiff-NCA model. The best results are achieved with our basic FourierDiff-NCA setup that uses 96 channels, a hidden size of 512, and 20 steps in the image space. Increasing the number of steps to $s=30$ negatively impacts the FID and KID values, as learning a meaningful rule becomes difficult. On the other hand, when $s=10$, the model does not have the perceptive range to perform proper diffusion in the image space. Decreasing the hidden size to $h=256$ or the channel size to $c=48$ both has a negative effect on FID and KID, emphasizing the need for a sufficiently large number of parameters to capture the distribution of the underlying dataset. The impact of a smaller channel size is less severe than the reduction of the hidden size, even though FourierDiff-NCA with $h=256$ has twice as many parameters as the setup with $c=48$. The results show that careful calibration of these parameters can lead to optimal performance, and fine-tuning this balance will be a focus of future work.
@@
Out-Of-Distribution Scale: Image Synthesis
Neural Cellular Automata have demonstrated the ability to generalize well beyond the specific training setup. Thus, we hypothesize that NCAs can handle larger image sizes and capture information beyond the pixel details of the input image through effective generalization. We synthesize images several times larger than the training size to test this. The results, shown in Figure 5, clearly show the ability of the model to not only produce face-like images at scales up to $128\times 128$ but also to add detail to the newly available space that is not present in the lower-resolution training images. This property is particularly evident in the finer details of the eyes and general facial features, which are more complex than observed at the training scale. As we expand the world’s scale to $78\times 64$, its content stretches while maintaining its relative orientation unchanged. Although some shortcomings in visual fidelity are apparent, it is likely that future improvements in the overall generation capabilities or OOD-specific optimizations of FourierDiff-NCA will improve OOD results.
Figure 6: FourierDiff-NCA can generate multiple predictions for a partially noisy image. The top row shows the input, while the other three rows show the predictions of the model.
Inpainting
@@
Figure 7: FourierDiff-NCA captures sub-pixel information during training. Due to the unique model architecture, we can upscale images beyond the training size of $64\times 64$ while adding additional details. In this case, we upscale $64\times 64$ real and synthetic images to $196\times 196$ images.
Image Upscaling
@@
Our investigation of FourierDiff-NCA showed that it can capture subpixel information and generate out-of-distribution (OOD) results. This intriguing result inspired us to explore the capacity of FourierDiff-NCA further. In this experiment, we investigate how well FourierDiff-NCA can upscale existing images to higher resolutions. To achieve this, we utilize real and synthetic low-resolution images of size $64\times 64$ and upscale them naively (using bilinear interpolation) to the desired image size of $192\times 192$. Then we add 10% noise, which we subsequently remove by utilizing FourierDiff-NCA. Notably, as we do not start from a fully noised image, we perform the backward diffusion for 10% of the training steps. The results are illustrated in Figure 7. Interestingly FourierDiff-NCA’s capability of upscaling $64\times 64$ synthesized images to $192\times 192$ is better than synthesizing the bigger scale from scratch. The model’s enhancement of details in the upscaled images confirms our original hypothesis that NCAs can generalize beyond the training size.
@@
Figure 8: Diff-NCA can train on images with a fixed resolution of $64\times 64$ pixels and infer different resolutions, e.g., $576\times 576$ pixels (top). The ability of Diff-NCA to generate different tissue types is shown in the bottom image, where eight generated samples are shown.
Fullscale Pathology: Image Synthesis
@@
A prevalent limitation of current diffusion and image generation models is that they are bound to fixed input sizes. This constraint becomes problematic when working with scenarios that require unconstrained world scales, such as digital pathology scans or satellite imagery. In such cases, global information is not essential to create an image that makes sense, as a coherent image can be synthesized by local communication alone. While UNet-based DDMs can generate individual patches quickly, synthesizing a seamless, continuous image that appears unified is beyond the capabilities of current architectures.
@@
Diff-NCA is designed to address these local constraints, allowing the generation of image patches of theoretically unlimited size while ensuring a seamless result. This is ensured by choosing a step size small enough that the model merely diffuses based on local features. In this process, local communication preserves the continuity of the image so that all components are aligned without disjointness or abrupt transitions. In Figure 8, we illustrate this by synthesizing a $576\times 576$ digital pathology scan, demonstrating the model’s ability to produce a seamless and visually coherent result. A key difference between increasing the world scale of Diff-NCA and FourierDiff-NCA after training is that Diff-NCA produces more tissue of the same scale, while FourierDiff-NCA increases its detail. Each of these methods holds its distinct utility and application. Diff-NCA’s ability to recognize and synthesize different tissue types and stains highlights its versatile and promising potential in unconstrained imaging tasks.
Conclusion
@@
We introduce Diff-NCA and FourierDiff-NCA as a jumpstart for Neural Cellular Automata-based denoising methods. Diff-NCA and FourierDiff-NCA are designed to produce high-quality images with merely 208k and 887k parameter counts, respectively. Diff-NCA is strategically optimized to utilize local features in specific applications such as pathology where they are sufficient. As an advancement of Diff-NCA, FourierDiff-NCA addresses the complex challenge of global communication within NCA by using the Fourier space for instantaneous global communication that is subsequently refined in image space. The effectiveness of these proposed architectures is demonstrated by generating $576\times 576$ pathology patches with Diff-NCA and $64\times 64$ CelebA images with FourierDiff-NCA. FourierDiff-NCA outperforms the NCA-based image generation model VNCA and five times bigger UNet-based DDMs. FourierDiff-NCA can perform super-resolution, out-of-distribution image generation, and inpainting without additional training. These results highlight the potential of Diff-NCA and FourierDiff-NCA as compelling alternative methods for efficient and versatile imaging that offer giga-pixel image generation in low-resource environments, democratizing generative learning.

### v2

@@ 4 Experimental Results
4.1 Data and Infrastructure
@@
CelebA dataset: The CelebA dataset (Liu et al., 2015), a widely used benchmark, consists of 202,599 images, each of size $178\times 218$. We scale all images to a uniform size of $64\times 64$ to match the dataset with the input requirements of the UNet and VNCA. The data is split 80%:10%:10% for training, validation, and testing. CelebA presents a challenge as it includes various facial images against different backgrounds. The inclusion of individuals with a range of accessories, hairstyles, and facial expressions further increases the complexity of the dataset. The size and variety of this dataset allow us to evaluate the ability of our models to handle intricate details and different visual features.
@@
BCSS Pathology dataset: This dataset (Amgad et al., 2019) contains 144 high-resolution pathology samples, which present a unique challenge for generative models. Initially, some images contain blur at the lowest level and are resized by a factor of four in the x and y direction to obtain clear and concise visual patterns. We then extract patches of size $64\times 64$ for training purposes. The BCSS pathology dataset provides an opportunity to rigorously test the capabilities of our proposed methods in generating large images in the medical domain. For this dataset, we use an 80%:10%:10% split for training, validation, and testing.
@@
Infrastructure: All models are implemented in PyTorch (Paszke et al., 2019), where we use the official implementation of VNCA (Palm et al., 2022). The models are trained on an Ubuntu 22.04 system using an Nvidia RTX 3090 and an Intel i7-12700 processor Additional details can be found in the appendix and the codebase that we will make available upon acceptance.
4.2 Metrics
@@
To assess the quality of the synthesized outputs, we use the well-established metrics Fréchet Inception Distance (FID) (Heusel et al., 2017) and Kernel Inception Distance (KID) (Bińkowski et al., 2018), which measure the similarity between the real and synthetic images based on the Inception-v3 (Szegedy et al., 2015) model. For both evaluations, we use a set of 2048 real images from the test split compared to an equal number of synthesized images.
4.3 Qualitative Comparison: Image Synthesis
@@
Examining the images, FourierDiff-NCA (1.85m) generated in Figure 3 shows that the model learns global information. Interestingly, FourierDiff-NCA does not have enough steps in the image space to achieve global knowledge transfer. Nevertheless, it exhibits such behavior, indicating that communication occurs in the Fourier space. Thanks to this Fourier communication, the model can capture various features of the underlying distribution, such as variations in facial features, hair color, clothing, and facial expressions, all while using a comparably small parameter count of 1.85m. The efficiency becomes apparent when comparing it to the four times bigger VNCA and the 2.5 times bigger UNet M-based DDM, where the results look worse with blurry results and substantial color shifts. Although the results of FourierDiff-NCA are imperfect, they represent a promising step toward detailed image generation with Neural Cellular Automata and demonstrate the potential for complex generative tasks.
4.4 Quantitative Comparison: Image Synthesis
@@
Figure 4 compares FourierDiff-NCA, various UNet configurations, and VNCA. Notably, FourierDiff-NCA excels at the tradeoff between parameters and FID reaching an FID of 43.86 and KID of 0.022 with only 1.1m parameters, outperforming four times larger UNet-based DDMs with an FID of 128.2 and KID of 0.089 and the ten times bigger VNCA with an FID of 299.99 and KID of 0.338. Exploring scalability, we train an enlarged FourierDiff-NCA variant with $c=128$ and $h=640$, dubbed FourierDiff-NCA (1.85m), 68% larger than the original. This model achieved a modest FID improvement at 42.69, but significantly enhanced the KID score to 0.016, indicating promising potential for future scaling of this method. This is further supported by the qualitative comparison in the appendix in Figure 13. In this figure, it is apparent, that while the 1.1m variant is struggling with general coherence, the 1.85m variant achieves better results in this regard. The notable improvements observed with the 1.85m parameter variant underscore its potential as a foundation for future research, encouraging further exploration into optimal configurations and scalability of the method. Although a large UNet with 111 million parameters produces the best results, performance drops with fewer parameters. Reducing the number of parameters in UNet configurations leads to a noticeable degradation in FID and KID metrics. The smallest UNet configuration with 652k parameters records an FID of 175.3 and a KID of 0.142. In contrast, FourierDiff-NCA, with 526k parameters, achieves an FID of 60.96 and a KID of 0.031. FourierDiff-NCA’s localized interaction architecture contributes to stable performance and improved parameter efficiency.
@@
Figure 5: Out-Of-Distribution image synthesis with FourierDiff-NCA (1.85m) of different scales and shapes.
4.5 Ablation
@@
The ablation study shown in Table 1 provides insights into the specific effects of the different hyperparameters of the FourierDiff-NCA model. The best results are achieved with our basic FourierDiff-NCA setup that uses 96 channels, a hidden size of 512, and 20 steps in the image space. Increasing the number of steps to $s=30$ negatively impacts the FID and KID values, as learning a meaningful rule becomes difficult. On the other hand, when $s=10$, the model does not have the perceptive range to incorporate enough local information to perform proper diffusion in the image space. Decreasing the hidden size to $h=256$ or the channel size to $c=48$ both has a negative effect on FID and KID, emphasizing the need for a sufficiently large number of parameters to capture the distribution of the underlying dataset. The impact of a smaller channel size is less severe than the reduction of the hidden size, even though FourierDiff-NCA with $h=256$ has twice as many parameters as the setup with $c=48$. The results show that careful calibration of these parameters can lead to optimal performance, and fine-tuning this balance will be a focus of future work.
@@
4.6 Out-Of-Distribution Scale: Image Synthesis
Neural Cellular Automata have demonstrated the ability to generalize well beyond the specific training setup. Thus, we hypothesize that NCAs can handle larger image sizes and capture information beyond the pixel details of the input image through effective generalization. We synthesize images several times larger than the training size to test this. The results, shown in Figure 5, clearly show the ability of the model to not only produce face-like images at scales up to $96\times 96$ but also to add detail to the newly available space that is not present in the lower-resolution training images. This property is particularly evident in the finer details of the eyes and general facial features, which are more complex than observed at the training scale. As we expand the world’s scale to $78\times 64$, its content stretches while maintaining its relative orientation unchanged. Although some shortcomings in visual fidelity are apparent, it is likely that future improvements in the overall generation capabilities or OOD-specific optimizations of FourierDiff-NCA will improve OOD results.
Figure 6: FourierDiff-NCA (1.85m) can generate multiple predictions for a partially noisy image. The top row shows the input, while the other three rows show the predictions of the model.
4.7 Inpainting
@@
Figure 7: FourierDiff-NCA (1.85m) captures sub-pixel information during training. Due to the unique model architecture, we can upscale images beyond the training size of $64\times 64$ while adding additional details. In this case, we upscale $64\times 64$ real and synthetic images to $128\times 128$ images.
4.8 Image Upscaling
@@
Our investigation of FourierDiff-NCA showed that it can capture subpixel information and generate out-of-distribution (OOD) results. This intriguing result inspired us to explore the capacity of FourierDiff-NCA further. In this experiment, we investigate how well FourierDiff-NCA can upscale existing images to higher resolutions. To achieve our objective, we start with real and synthetic low-resolution images of size $64\times 64$. These images are initially upscaled to $128\times 128$ using a naive approach with nearest-neighbor interpolation. In this upscaling process, we essentially create a grid where the original pixels form the ’1’ positions in a pattern of [[1, 0], [0, 0]], leaving the ’0’ positions as newly added pixels. After this upscaling, we introduce 90% noise and then employ FourierDiff-NCA for denoising. Importantly, in the backward diffusion process, we specifically target only the newly added pixels for updates. Additionally, during each update, we blend in 2% of the original image data to these pixels, enhancing the integration of the upscaled and original image components. The results are illustrated in Figure 7. FourierDiff-NCA not only introduces new details in upscaled images but also integrates these with existing information, enhancing overall detail for a richer high-resolution output, affirming our original hypothesis that NCAs can generalize beyond the training size.
@@
Figure 8: Diff-NCA can train on images with a fixed resolution of $64\times 64$ pixels and infer different resolutions, e.g., $512\times 512$ pixels (top). The ability of Diff-NCA to generate different tissue types is shown in the bottom image, where eight generated samples are shown.
4.9 Fullscale Pathology: Image Synthesis
@@
A prevalent limitation of current diffusion and image generation models is that they are bound to fixed input sizes. This constraint becomes problematic when working with scenarios that require unconstrained world scales, such as digital pathology scans or satellite imagery. In such cases, global information is not essential to create an image that makes sense, as a coherent image can be synthesized by local communication alone. While UNet-based DDMs can generate individual patches, synthesizing a seamless, continuous image that appears unified is beyond the capabilities of current architectures.
@@
Diff-NCA is designed to address these local constraints, allowing the generation of image patches of theoretically unlimited size while ensuring a seamless result. This is ensured by choosing a step size small enough that the model merely diffuses based on local features. In this process, local communication preserves the continuity of the image so that all components are aligned without disjointness or abrupt transitions. In Figure 8, we illustrate this by synthesizing a $512\times 512$ digital pathology scan, demonstrating the model’s ability to produce a seamless and visually coherent result. A key difference between increasing the world scale of Diff-NCA and FourierDiff-NCA after training is that Diff-NCA produces more tissue of the same scale, while FourierDiff-NCA increases its detail since it has been trained on integrating the global information. Each of these methods holds its distinct utility and application. Diff-NCA’s ability to recognize and synthesize different tissue types and stains highlights its versatile and promising potential in unconstrained imaging tasks.
5 Conclusion
@@
We introduce Diff-NCA and FourierDiff-NCA as a jumpstart for Neural Cellular Automata-based denoising methods. Diff-NCA and FourierDiff-NCA are designed to produce high-quality images with merely 331k and 1.1m parameter counts, respectively. Diff-NCA is strategically optimized to utilize local features in specific applications such as pathology where they are sufficient. As an advancement of Diff-NCA, FourierDiff-NCA addresses the complex challenge of global communication within NCA by using the Fourier space for instantaneous global communication that is subsequently refined in image space. The effectiveness of these proposed architectures is demonstrated by generating $512\times 512$ pathology patches with Diff-NCA and $64\times 64$ CelebA images with FourierDiff-NCA. FourierDiff-NCA outperforms the NCA-based image generation model VNCA and four times bigger UNet-based DDMs. FourierDiff-NCA can perform super-resolution, out-of-distribution image generation, and inpainting without additional training. These results highlight the potential of Diff-NCA and FourierDiff-NCA as compelling alternative methods for efficient and versatile imaging that offer giga-pixel image generation in low-resource environments, democratizing generative learning.
Impact Statement
In this paper, we introduce Diff-NCA and FourierDiff-NCA, Neural Cellular Automata-based DDM architectures that significantly reduce parameter count in image synthesis, compared to conventional UNet architectures. This reduction presents a notable advancement in model compactness, potentially leading to broader applications, especially in environments with limited computational resources.
Appendix A Neural Cellular Automata (NCA): Introduction
Neural Cellular Automata (NCA) represent a type of minimal model that deviates from traditional deep learning architectures by focusing on individual pixels or cells within an image. This departure from a holistic view of the entire scene contributes to their compactness. This means that a single NCA looks only at one cell. To take into account global information, this same rule is replicated across all pixels of an image. Recognizing the need for information exchange beyond their individual states, these models are fed by an input layer-often in the form of a 3x3 convolutional layer to establish cellular communication with neighboring cells. A single update step of one cell is illustrated in Figure 9. Consequently, each cell acquires awareness of its own state while engaging in controlled interactions with its immediate environment. To handle more complicated and multi-faceted tasks, the NCA approach requires an iterative execution that allows the model to operate over multiple temporal iterations. This iterative approach facilitates the sharing of information across the entirety of the image, leveraging the ability to successfully perform complex operations orchestrated on a broader spatial scale.
Figure 9: An update step of a single cell.
In summary, Neural Cellular Automata represent an efficient architecture in which updates occur at the level of individual cells that are equipped with neighborhood cell communication. This approach, combined with the iterative application of said update rule, enables these models to handle complicated image analysis problems by orchestrating a coordinated flow of information across the entire image.
A.1 Important concepts
When considering the technical implementation of Neural Cellular Automata (NCAs), several key concepts need to be explored. First and foremost is the step count, which indicates the number of times the model is applied at the image scale. This parameter largely determines the perceptive range of a single NCA model and limits the communicative range. The perceptive range is further influenced by the fire rate that controls the stochastic cell updates, implemented by a stochastic reset mechanism. Furthermore, the firing rate serves as a critical factor in shaping a robust update rule within the model and enhances its ability to develop a more robust and adaptive rule.
A.1.1 VRAM requirements
Despite their lightweight architecture, training NCAs can impose significant demands on VRAM as the learned update rule is replicated across the whole image scale. Combined with the iterative model execution, this leads to an exponential increase in memory requirements highly dependent on image size. As all states have to be kept in the VRAM to perform backpropagation via gradient decent, a naive approach to running an NCA on a $256\times 256$ image for only 60 steps results in VRAM usage of over 20 GB (Kalkhof et al., 2023).
Figure 10: Increase of the perceptive range of the red cell in image space.
A.1.2 Perceptive Range / Communication
The perceptive range of an NCA increases with its number of steps. In a naive setup, which uses a 3x3 convolution as the input to the model, each NCA can only communicate with its direct neighbors. This means propagating information across a 100x100 image subsequentially requires 100 steps. This makes it difficult and slow to acquire global information. Further, it has to be carefully balanced, as a perceptive range too big, makes it difficult for the model to learn a meaningful rule, while one that is too small, prevents the model from gaining global insight. It is further influenced by the fire rate, as a stochastic reset implies that this cell won’t receive any updates in this timestep.
A.1.3 Fire Rate
In the context of NCAs, a common concept is the introduction of a firing rate that controls cellular updates. This rate, which is typically 50%, determines whether the state of a cell changes in each iteration. The fire rate is implemented by canceling the cellular update for the corresponding step, by applying a stochastic reset. Consequently, the model is forced to adapt to the stochastic flow of information, which increases its robustness, as shown in the growth of cells from singular pixels (Mordvintsev et al., 2020).
A.1.4 Channels
The channels used include both input data, such as the image or in our case the noisy image, and additional channels that are used for storing information inbetween the NCAs update steps. Ensuring adequate channel size is critical, as an insufficient capacity can compromise the ability of the NCA to store important information between steps.
Appendix B Fourier Communication
The Fourier space is characterized by a natural arrangement in which the low-frequency information is concentrated in the center of the space. Notably, a significant portion of the important low-frequency data is preserved even when 93.75% of the information is removed, as illustrated in Figure 11. This persistence of essential low-frequency content has sufficient information value to allow orientation in a global context. The unique property of Fourier space to retain key information despite significant data reduction underscores its suitability for retaining fundamental spatial features across scales.
Figure 11: Cutting away 93.75% of the information in Fourier space drastically reduces the size, yet important low-frequency information is preserved.
Figure 12: Simplified visualization, of how communication between NCA’s flows in the Fourier space and how it translates to the image space. A single step is enough to communicate across the whole image, whereas extra steps refine the signal sent.
Taking advantage of its inherent structural features, the Fourier space turns out to be a strategically advantageous channel for accelerating global communication within local interaction models. Said process involves the NCA performing a particular number of steps in Fourier space, leading to the accumulation of comprehensive global knowledge. Following this phase, the accumulated knowledge is transferred into the image space through an inverse Fast Fourier Transform. Figure 12 provides a simplified representation of this information flow, illustrating the interplay between Fourier space and image space. This novel use of the characteristics of Fourier space highlights its central role in enabling efficient global communication of locally interacting NCAs.
Appendix C Qualitative Comparison FourierDiff-NCAs
A qualitative comparison of FourierDiff-NCA models with 1.1m and 1.85m parameters is presented in Figure 13. Despite a modest 3% increase in FID, the KID score sees a substantial 28% improvement, reflecting the enhanced coherence of the generated images.
Figure 13: Qualitative comparison between FourierDiff-NCA 1.1m and FourierDiff-NCA 1.85m.
Appendix D Quantitative Results
The precise quantitative results comparing FourierDiff-NCA, VNCA, and the UNets (S, M, B, L) compared in our figure can be found in Table 2.
Method FID $\downarrow$ KID $\downarrow$ # Param. $\downarrow$
FD-NCA (1.85m) 42.69 0.016 $\pm$ 0.012 1,849,632
FD-NCA 43.86 0.022 $\pm$ 0.015 1,101,216
VNCA 299.99 0.338 $\pm$ 0.032 9,732,416
UNet S 175.3 0.142 $\pm$ 0.009 652,195
UNet M 128.2 0.089 $\pm$ 0.009 3,940,547
UNet B 20.1 0.005 $\pm$ 0.002 17,217,923
UNet L 18.2 0.004 $\pm$ 0.002 111,051,779
Table 2: Quantiative comparison of FourierDiff-NCA (here FD-NCA) and UNet-based DDMs.
Appendix E Experimental Setup
To ensure better reproducibility, detailed experiment settings are given below. All experiments performed are done with PyTorch (Paszke et al., 2019), with comprehensive descriptions complementing the full codebase, which will be released after acceptance.
UNet Channel dimension Resnet block per resolution Attention resolution $\downarrow$ # Param.
Small (32; 64) 1 (-) 652,195
Medium (32; 64; 128) 2 (16) 3,940,547
Big (64; 128; 128; 256) 2 (16) 17,217,923
Large (128; 128; 256; 256; 512; 512) 2 (16; 8) 111,051,779
Table 3: Configuration of UNets.
E.1 Diff-NCA and FourierDiff-NCA
All conducted experiments utilize the Adam optimizer (Kingma & Ba, 2014). The chosen hyperparameters include a learning rate of $1.6\times 10^{-3}$, a learning rate gamma of $0.9999$, betas for the learning rate as $(0.9,0.99)$, and epsilon ($\epsilon$) set at $1\times 10^{-8}$. The models undergo training for 200,000 steps, utilizing a batch size of 16. Detailed configurations for FourierDiff-NCA are outlined in Listing LABEL:lst:FourierDiff, while those for Diff-NCA are provided in Listing LABEL:lst:Diff.
FourierDiff_NCA(
(norm_four): GroupNorm(1, 512, eps=1e-05, affine=True)
(perceive_four): Conv2d(196, 192, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
(fc0_four): Conv2d(388, 512, kernel_size=(1, 1), stride=(1, 1))
(fc1_four): Conv2d(512, 192, kernel_size=(1, 1), stride=(1, 1))
(mulCond0_four): Sequential(
(0): Conv2d(4, 128, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(128, 388, kernel_size=(1, 1), stride=(1, 1))
(mulCond1_four): Sequential(
(0): Conv2d(4, 128, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(128, 512, kernel_size=(1, 1), stride=(1, 1))
(embedding_four): Sequential(
(0): Conv2d(16, 256, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(256, 4, kernel_size=(1, 1), stride=(1, 1))
(norm_img): GroupNorm(1, 512, eps=1e-05, affine=True)
(perceive_img): Conv2d(100, 96, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), padding_mode=reflect)
(fc0_img): Conv2d(196, 512, kernel_size=(1, 1), stride=(1, 1))
(fc1_img): Conv2d(512, 96, kernel_size=(1, 1), stride=(1, 1))
(mulCond0_img): Sequential(
(0): Conv2d(4, 128, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(128, 196, kernel_size=(1, 1), stride=(1, 1))
(mulCond1_img): Sequential(
(0): Conv2d(4, 128, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(128, 512, kernel_size=(1, 1), stride=(1, 1))
(embedding_img): Sequential(
(0): Conv2d(16, 256, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(256, 4, kernel_size=(1, 1), stride=(1, 1))
Diff_NCA(
(norm_img): GroupNorm(1, 512, eps=1e-05, affine=True)
(perceive_img): Conv2d(98, 96, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1), padding_mode=reflect)
(fc0_img): Conv2d(194, 512, kernel_size=(1, 1), stride=(1, 1))
(fc1_img): Conv2d(512, 96, kernel_size=(1, 1), stride=(1, 1))
(mulCond0_img): Sequential(
(0): Conv2d(2, 128, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(128, 194, kernel_size=(1, 1), stride=(1, 1))
(mulCond1_img): Sequential(
(0): Conv2d(2, 128, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(128, 512, kernel_size=(1, 1), stride=(1, 1))
(embedding_img): Sequential(
(0): Conv2d(4, 256, kernel_size=(1, 1), stride=(1, 1))
(1): SiLU()
(2): Conv2d(256, 2, kernel_size=(1, 1), stride=(1, 1))
E.2 VNCA
We use the official CelebA specific implementation of VNCA (Palm et al., 2022) which can be found at https://github.com/rasmusbergpalm/vnca.
E.3 UNets
The training of all UNets employs the Adam optimizer with a learning rate of $3\times 10^{-5}$, betas set at $(0.5,0.999)$, and epsilon ($\epsilon$) at $1\times 10^{-6}$. A training duration of 200,000 steps is undertaken using a batch size of 32. Refer to Table 3 for detailed layer configurations of each UNet model.